OpenStudy (anonymous):

why is tan 90 undefined

terenzreignz (terenzreignz):

Well, try the known identity \[\Large \tan(\theta) = \frac{\sin(\theta)}{\cos(\theta)}\]

OpenStudy (anonymous):

what does this mean?

OpenStudy (unklerhaukus):

a triangle can only have one right angle

terenzreignz (terenzreignz):

Try substituting \(\large \theta = 90 ^o\)

OpenStudy (anonymous):

Is it because it is perpendicular?

terenzreignz (terenzreignz):

Not so hot on why....sohcahtoa?

OpenStudy (amistre64):

tangent can be viewed as the slope of a line y/x  as the angle \(\alpha\) gets larger and larger, it approaches 90 degrees. the slope of the line, it we let x=1, increases towards infinity, which is not a definable value.